# EqNIO: Subequivariant Neural Inertial Odometry

### v1

@@ Abstract
Presently, neural networks are widely employed to accurately estimate 2D displacements and associated uncertainties from Inertial Measurement Unit (IMU) data that can be integrated into stochastic filter networks like the Extended Kalman Filter (EKF) as measurements and uncertainties for the update step in the filter. However, such neural approaches overlook symmetry which is a crucial inductive bias for model generalization. This oversight is notable because (i) physical laws adhere to symmetry principles when considering the gravity axis, meaning there exists the same transformation for both the physical entity and the resulting trajectory, and (ii) displacements should remain equivariant to frame transformations when the inertial frame changes. To address this, we propose a subequivariant framework by: (i) deriving fundamental layers such as linear and nonlinear layers for a subequivariant network, designed to handle sequences of vectors and scalars, (ii) employing the subequivariant network to predict an equivariant frame for the sequence of inertial measurements. This predicted frame can then be utilized for extracting invariant features through projection, which are integrated with arbitrary network architectures, (iii) transforming the invariant output by frame transformation to obtain equivariant displacements and covariances. We demonstrate the effectiveness and generalization of our Equivariant Framework on a filter-based approach with TLIO architecture for TLIO and Aria datasets, and an end-to-end deep learning approach with RONIN architecture for RONIN, RIDI and OxIOD datasets.
@@ 3 Problem Setup
This paper aims to estimate the orientation, position, and/or velocity of a rigid body over time given a sequence of IMU accelerometer (giving accelerations $a_{i}\in\mathbb{R}^{3}$) and gyroscope (giving angular velocity $\omega_{i}\in\mathbb{R}^{3}$) measurements $\{\{a_{i},\omega_{i}\}\}_{i=1}^{n}$, expressed in the local IMU inertial frame. Simply integrating these measurements results in large drift due to sensor noise and biased IMU readings, and thus, we use off-the-shelf neural networks to regress accurate 3 degrees-of-freedom (DoF) linear velocities [30] or 3D displacement measurements $d\in\mathbb{R}^{3}$ and covariances $\Sigma\in\mathbb{R}^{3\times 3}$ [33]. In the latter case, we treat the network outputs as measurements and fuse them in an Extended Kalman Filter (EKF) estimating the IMU state, i.e. orientation, position, velocity, IMU biases, and uncertainties. Preliminaries on the terms used in inertial odometry are included in Appendix A.1.2. Preliminaries and details on EKFs and the associated IMU measurement model are included in Appendix A.4.
@@
To reduce the data variability in IMU measurements, we map them to a gravity-aligned frame (i.e. perform gravity-alignment), which has its z-axis aligned with the gravity direction. We do this by simple rotations. Typically, the gravity direction can be estimated from an auxiliary visual-inertial odometry (VIO) method or can be estimated from the dominant direction of accelerometer readings at rest. Reducing the data variability aids in learning priors associated with particular IMU motions, and also simplifies odometry due to the elimination of two degrees of freedom. This frame is, however, ill-defined, since simply rotating it around z or reflecting it across planes parallel to z (applications of rotations or roto-reflections from the groups $SO(2)$ and $O(2)$) result in new valid gravity-aligned frames. Our main contribution is to eliminate these additional degrees of variability in IMU data with an $\bm{SO(2)}$ or $\bm{O(2)}$ equivariance framework, and thus facilitate more efficient learning of motion priors. Our framework operates in three steps as seen in Figure 2(a): After gravity alignment, it first estimates a canonical frame using an equivariant neural network. Then, it maps the IMU data into this frame, i.e. canonicalizes them, before predicting linear velocity or displacement and covariance with an off-the-shelf neural network $\Phi$. Finally, it uses the canonical frame to map the resulting quantities into the original frame. In the language of equivariance research, our method thus has the property that transforming the gravity-aligned frame with the aforementioned symmetries (and as a result, the IMU data expressed in that frame) leads to an equivariant change in the resulting displacement and covariances. Unlike related approaches [30, 33, 8], which enforce this property with yaw augmentations or auxiliary losses (i.e. possess non-strict equivariance), our method enforces this property exactly (strict $O(2)$ equivariance), and by design.
Denote $R_{3\times 3}\in\mathbb{R}^{3\times 3}\subset O(3)$ an orthonormal roto-reflection operator which rotates vectors around the $z$-axis, and reflects them along a plane parallel to the $z$-axis if $\text{det}(R)=-1$. Figure 3 illustrates the effect of $R_{3\times 3}$ on displacements and the resulting trajectory. Since $R$ only affects the $xy$ components of vectors $v$ and keeps their $z$ components, we decompose $R_{3\times 3}=R_{2\times 2}\oplus 1$ (where $\oplus$ composes $R_{2\times 2}$ and $1$ in block diagonal form), and $v=[v_{xy}\,v_{z}]$ into their separate components. Since $R_{3\times 3}$ can be decomposed, we identify $R_{3\times 3}$ with $R_{2\times 2}$ and thus treat the reduced $SO(2)$ or $O(2)$ subequivariance problem. We formalize and achieve equivariance with the tools of group and representation theory, and include necessary preliminaries in the Appendix A.1.1. We train a network $\Phi$ to regress displacements and potentially covariances from IMU measurements, and thus
$d,\Sigma=\Phi(\{a_{i},\omega_{i}\}_{i=1}^{n})$ (1)
@@
We design separate networks that handle $SO(2)$ and $O(2)$ subequivariance, i.e. have the property $\Phi(\{\rho_{a}(R_{2\times 2})a_{i},\rho_{\omega}(R_{2\times 2})\omega_{i}\}_{i=1}^{n})=\rho_{o}(R_{2\times 2})\Phi(\{a_{i},\omega_{i}\}_{i=1}^{n})$. Here $\rho$ denotes a group representation and generates suitable transformation matrices for acceleration $(\rho_{a})$, angular velocity $(\rho_{\omega})$, and output variables $\rho_{o}$ under the action of $R_{2\times 2}$. The representation $\rho_{o}$ in particular decomposes into $\rho_{o}=\rho_{d}\oplus\rho_{\Sigma}$, where $\rho_{d}$ and $\rho_{\Sigma}$ act separately on the displacement and covariance generated by the neural network. For accelerations and displacements, $\rho_{a}(R_{2\times 2})=R_{2\times 2}\oplus 1$ and $\rho_{d}(R_{2\times 2})=R_{2\times 2}\oplus 1$, while for angular velocities $\rho_{\omega}(R_{2\times 2})=\text{det}(R_{2\times 2})(R_{2\times 2}\oplus 1)$. Upon reflection, the rotation direction of $\omega$ must be additionally inverted, which is why we add a factor of $\text{det}(R)$, and is further illustrated in Figure 4(a). When designing the equivariant network, we desire all input quantities to transform similarly. For this reason, when dealing with $O(2)$ equivariance, we will later introduce an additional preprocessing step, which decomposes $\omega=v_{1}\times v_{2}$. These vectors transform in the same way as acceleration, i.e. $\rho_{v_{2}}(R_{2\times 2})=\rho_{v_{1}}(R_{2\times 2})=R_{2\times 2}\oplus 1$, and lead to the desired transform of resulting $\omega$. Finally, covariance $\Sigma$ transforms as $(R_{2\times 2}\oplus 1)\Sigma(R_{2\times 2}\oplus 1)^{T}$. If we write $\Sigma$ in vectorized form we have $\text{vec}(\Sigma)^{\prime}=(R_{2\times 2}\oplus 1)\otimes(R_{2\times 2}\oplus 1)\text{vec}(\Sigma)$ so $\rho_{\Sigma}(R_{2\times 2})=(R_{2\times 2}\oplus 1)\otimes(R_{2\times 2}\oplus 1)$. Here $\otimes$ denotes the Kronecker product between matrices.
4 Equivariant Framework
@@ 4.1 Preprocessing
As mentioned in Section 3, accelerations and angular velocities do not transform the same way under an action of $R_{2\times 2}\in O(2)$ when there is a reflection involved. This means that we cannot apply the same $O(2)$ equivariant network for both IMU data components. To address this, we decompose the angular rate into vectors $v_{1},v_{2}$ such that $\omega=v_{1}\times v_{2}$ (see Figure 4(b)). Both vectors transform as $v^{\prime}_{1/2}=(R_{2\times 2}\oplus 1)v_{1/2}$ under rotation $R_{2\times 2}$, identical to the acceleration, and their cross product has the desirable property $\omega^{\prime}=v^{\prime}_{1}\times v^{\prime}_{2}=((R_{2\times 2}\oplus 1)v_{1})\times((R_{2\times 2}\oplus 1)v_{2})=\text{det}(R_{2\times 2}\oplus 1)(R_{2\times 2}\oplus 1)(v_{1}\times v_{2})=\text{det}(R_{2\times 2})(R_{2\times 2}\oplus 1)\omega$, using the standard cross-product property $(Ax)\times(Ay)=\text{det}(A)A(x\times y)$. The group action on $\omega$ exactly coincides with what was derived in Section 3. Our preprocessing step is summarized by the following mapping $\mathcal{F}:\mathbb{R}^{6}\rightarrow\mathbb{R}^{9}$ defined as:
@@
[…]
@@ 4.2 Equivariant Frame
Instead of enforcing the above equivariance constraint on every network layer, we instead employ an equivariant network (see Figure 2(b)) to regress a canonical frame $G$ from the preprocessed input obtained from preprocessed IMU measurements $\{a_{i},\omega_{i}\}_{i=1}^{n}$ when dealing with $SO(2)$ and preprocessed IMU measurements $\{a_{i},v_{1,i},v_{2,i}\}_{i=1}^{n}$, when dealing with $O(2)$. The canonical frame $G$ is uniquely achieved by transforming the identity frame with $R_{G}\in SO(2)$ or $O(2)$, which means we equivalently learn the transformation $R_{G}$ when we learn the equivariant frame $G$. We first convert the IMU data into $n\times C_{0}^{s}$ rotation invariant scalar features, and $n\times 2\times C_{0}^{v}$ rotation equivariant vector features, and process them jointly, taking inspiration from [44], which shows that learning from invariant features alongside vector inputs can produce universally equivariant outputs. As vector features we select the $xy$-components of each input vector ($C_{0}^{v}=2$ for $SO(2)$ and $C_{0}^{v}=3$ for $O(2)$). Instead, as scalar features we select (i) the z-components of each vector, (ii) the norm of the $xy$-components of each vector, and (iii) the pairwise dot-product of the $xy$-components of each vector. For $SO(3)$ we have $C_{0}^{s}=2+2+1=5$, while for $O(2)$ we have $C_{0}^{s}=3+3+3=9$.
As seen in Figure 2(a), preprocessed IMU measurements $\{a_{i},\omega_{i}\}_{i=1}^{n}$ are mapped into equivariant frame $G$ i.e. canonicalized into invariant features $\{a^{\prime}_{i},\omega^{\prime}_{i}\}_{i=1}^{n}$. This mapping is $a^{\prime}_{i}=\rho_{a}(R_{G}^{T})a_{i}$, and similarly for other vectors. Note that we first perform the aforementioned preprocessing when dealing with $O(2)$. We then regress the invariant displacement $d^{\prime}$ and covariance $\Sigma^{\prime}$ on these invariant features before mapping back into the original frame, using the same canonical frame. Specifically, the invariant displacement is remapped as $\text{vec}(\Sigma)=\rho_{\Sigma}(R_{G})\Sigma$. This results in equivariant displacement $d=\rho_{d}(R_{G})d^{\prime}$ and covariance $\text{vec}(\Sigma)=\rho_{\Sigma}(R_{G})\text{vec}(\Sigma^{\prime})$. This design enables seamless integration of our equivariant network while separating the effects of transformations, leveraging traditional networks to process intrinsic data.

### v2

@@ Abstract
Neural networks are seeing rapid adoption in purely inertial odometry, where accelerometer and gyroscope measurements from commodity inertial measurement units (IMU) are used to regress displacements and associated uncertainties. They can learn informative displacement priors, which can be directly fused with the raw data with off-the-shelf non-linear filters. Nevertheless, these networks do not consider the physical roto-reflective symmetries inherent in IMU data, leading to the need to memorize the same priors for every possible motion direction, which hinders generalization. In this work, we characterize these symmetries and show that the IMU data and the resulting displacement and covariance transform equivariantly, when rotated around the gravity vector and reflected with respect to arbitrary planes parallel to gravity. We design a neural network that respects these symmetries by design through equivariant processing in three steps: First, it estimates an equivariant gravity-aligned frame from equivariant vectors and invariant scalars derived from IMU data, leveraging expressive linear and non-linear layers tailored to commute with the underlying symmetry transformation. We then map the IMU data into this frame, thereby achieving an invariant canonicalization that can be directly used with off-the-shelf inertial odometry networks. Finally, we map these network outputs back into the original frame, thereby obtaining equivariant covariances and displacements. We demonstrate the generality of our framework by applying it to the filter-based approach based on TLIO, and the end-to-end RONIN architecture, and show better performance on the TLIO, Aria, RIDI and OxIOD datasets than existing methods.
@@ 3 Problem Setup
This paper aims to estimate the orientation, position, and/or velocity of a rigid body over time given a sequence of IMU accelerometer (giving accelerations $a_{i}\in\mathbb{R}^{3}$) and gyroscope (giving angular velocity $\omega_{i}\in\mathbb{R}^{3}$) measurements $\{(a_{i},\omega_{i})\}_{i=1}^{n}$, expressed in the local IMU inertial frame. Simply integrating these measurements results in large drift due to sensor noise and biased IMU readings, and thus, we use off-the-shelf neural networks to regress accurate 3 degrees-of-freedom (DoF) linear velocities [30] or 3D displacement measurements $d\in\mathbb{R}^{3}$ and covariances $\Sigma\in\mathbb{R}^{3\times 3}$ [33]. In the latter case, we treat the network outputs as measurements and fuse them in an Extended Kalman Filter (EKF) estimating the IMU state, i.e. orientation, position, velocity, IMU biases, and uncertainties. Preliminaries on the terms used in inertial odometry are included in Appendix A.1.2. Preliminaries and details on EKFs and the associated IMU measurement model are included in Appendix A.4.
@@
To reduce the data variability in IMU measurements, we map them to a gravity-aligned frame (i.e. perform gravity-alignment), which has its z-axis aligned with the gravity direction. We do this by simple rotations. Typically, the gravity direction can be estimated from an auxiliary visual-inertial odometry (VIO) method or can be estimated from the dominant direction of accelerometer readings at rest during training. During test time, the orientation estimated from the current state of the EKF or known orientation is used for gravity-alignment. Reducing the data variability aids in learning priors associated with particular IMU motions, and also simplifies odometry due to the elimination of two degrees of freedom. This frame is, however, ill-defined, since simply rotating it around z or reflecting it across planes parallel to z (applications of rotations or roto-reflections from the groups $SO(2)$ and $O(2)$) result in new valid gravity-aligned frames. Our main contribution is to eliminate these additional degrees of variability in IMU data with an $\bm{SO(2)}$ or $\bm{O(2)}$ equivariance framework, and thus facilitate more efficient learning of motion priors.
Our framework operates in three steps as seen in Figure 2(a): After gravity alignment, it first estimates a canonical frame using an equivariant neural network. Then, it maps the IMU data into this frame, i.e. canonicalizes them, before predicting linear velocity or displacement and covariance with an off-the-shelf neural network $\Phi$. Finally, it uses the canonical frame to map the resulting quantities into the original frame. In the language of equivariance research, our method thus has the property that transforming the gravity-aligned frame with the aforementioned symmetries (and as a result, the IMU data expressed in that frame) leads to an equivariant change in the resulting displacement and covariances. Unlike related approaches [30, 33, 8], which enforce this property with yaw augmentations or auxiliary losses (i.e. possess non-strict equivariance), our method enforces this property exactly (strict $O_{\vec{g}}(3)$ equivariance), and by design.
@@
Similar to prior works [15, 28], the presence of gravity breaks symmetry in the z-axis which is aligned with the gravity $\vec{g}\in\mathbb{R}^{3}$. Hence, the orthogonal symmetry $O(3)$ is no longer preserved in all directions. In fact, it is only preserved for the subgroup $O_{\vec{g}}(3):=\{R\in O(3)|R\vec{g}=\vec{g}\}$ where $R\in O(3):=\{R\in\mathbb{R}^{3\times 3}|R^{T}R=I\}$. This relaxation of strict $O(3)$ equivariance in the gravity direction is termed as $O_{\vec{g}}(3)$ equivariance or $O(3)$ subequivariance. Similarly, the subgroup $SO_{\vec{g}}(3):=\{R\in SO(3)|R\vec{g}=\vec{g}\}$ where $R\in SO(3):=\{R\in\mathbb{R}^{3\times 3}|R^{T}R=I,\text{det}(R)=1\}$. Denote $R_{3\times 3}\in\mathbb{R}^{3\times 3}\subset O_{\vec{g}}(3)$ an orthonormal roto-reflection operator which rotates vectors around the $z$-axis (axis aligned with gravity direction), and reflects them along a plane parallel to the $z$-axis if $\text{det}(R_{3\times 3})=-1$. Figure 3 illustrates the effect of $R_{3\times 3}$ on displacements and the resulting trajectory. Since $R_{3\times 3}$ only affects the $xy$ components of vectors $v$ and keeps their $z$ components, we decompose $R_{3\times 3}=R_{2\times 2}\oplus 1$ (where $\oplus$ composes $R_{2\times 2}$ and $1$ in block diagonal form), and $v=(v_{xy},v_{z})$ into their separate components. Since $R_{3\times 3}$ is fully defined by $R_{2\times 2}$, we can directly identify it with the rotation matrix $R_{2\times 2}$ which is part of the smaller subgroup $SO(2):=\{R\in\mathbb{R}^{2\times 2}|R^{T}R=I,\text{det}(R)=1\}$ or $O(2):=\{R\in\mathbb{R}^{2\times 2}|R^{T}R=I\}$. In fact, the above argument implies that $SO(2)$ and $SO_{\vec{g}}(3)$ are isomorphic, i.e. $SO(2)\cong SO_{\vec{g}}(3)$ (and similarly $O_{\vec{g}}(3)\cong O(2)$) and so for the remainder of the paper we will simply use the terms $SO(2)$ and $O(2)$ equivariance. Moreover, we refer to the xy-components $v_{xz}$ of vectors that transform with $R_{2\times 2}$ as equivariant vectors and the z-components $v_{z}$ that remain unchanged by $R_{2\times 2}$ as invariant scalars. We formalize and achieve equivariance with the tools of group and representation theory, and include necessary preliminaries in the Appendix A.1.1. We train an off-the-shelf neural network $\Phi:\mathbb{R}^{n\times 2\times 3}\rightarrow\mathbb{R}^{3}\times\ \mathbb{R}^{3\times 3}$ that takes as input a sequence of $n$ IMU measurements $\{(a_{i},\omega_{i})\}_{i=1}^{n}$ in a gravity-aligned frame and regresses 3D displacement ($d\in\mathbb{R}^{3}$) and covariance ($\Sigma\in\mathbb{R}^{3\times 3}$) in the same frame, and thus
$d,\Sigma=\Phi(\{(a_{i},\omega_{i})\}_{i=1}^{n})$ (1)
@@
We design separate networks that handle $SO(2)$ and $O(2)$ equivariance, i.e. have the property $\Phi(\{(\rho_{a}(R_{2\times 2})a_{i},\rho_{\omega}(R_{2\times 2})\omega_{i})\}_{i=1}^{n})=\rho_{o}(R_{2\times 2})\Phi(\{(a_{i},\omega_{i})\}_{i=1}^{n})$. Here $\rho$ denotes a group representation and generates suitable transformation matrices for acceleration $(\rho_{a})$, angular velocity $(\rho_{\omega})$, and output variables $\rho_{o}$ under the action of $R_{2\times 2}$. The representation $\rho_{o}$ in particular decomposes into $\rho_{o}=\rho_{d}\oplus\rho_{\Sigma}$, where $\rho_{d}$ and $\rho_{\Sigma}$ act separately on the displacement and covariance generated by the neural network. For accelerations and displacements, $\rho_{a}(R_{2\times 2})=R_{2\times 2}\oplus 1$ and $\rho_{d}(R_{2\times 2})=R_{2\times 2}\oplus 1$, while for angular velocities $\rho_{\omega}(R_{2\times 2})=\text{det}(R_{2\times 2})(R_{2\times 2}\oplus 1)$. Upon reflection, the rotation direction of $\omega$ must be additionally inverted, which is why we add a factor of $\text{det}(R)$, and is further illustrated in Figure 4(a). When designing the equivariant network, we desire all input quantities to transform similarly. For this reason, when dealing with $O_{\vec{g}}(3)$ equivariance, we will later introduce an additional preprocessing step, which decomposes $\omega=v_{1}\times v_{2}$. These vectors transform in the same way as acceleration, i.e. $\rho_{v_{2}}(R_{2\times 2})=\rho_{v_{1}}(R_{2\times 2})=R_{2\times 2}\oplus 1$, and lead to the desired transform of resulting $\omega$. Finally, covariance $\Sigma$ transforms as $(R_{2\times 2}\oplus 1)\Sigma(R_{2\times 2}\oplus 1)^{T}$. If we write $\Sigma$ in vectorized form we have $\text{vec}(\Sigma)^{\prime}=(R_{2\times 2}\oplus 1)\otimes(R_{2\times 2}\oplus 1)\text{vec}(\Sigma)$ so $\rho_{\Sigma}(R_{2\times 2})=(R_{2\times 2}\oplus 1)\otimes(R_{2\times 2}\oplus 1)$. Here $\otimes$ denotes the Kronecker product between matrices.
4 Subequivariant Framework
@@ 4.1 Preprocessing
As mentioned in Section 3, accelerations and angular velocities do not transform the same way under an action of $R_{2\times 2}\in O(2)$ when there is a reflection involved. To address this, we decompose the angular rate into vectors $v_{1},v_{2}$ such that $\omega=v_{1}\times v_{2}$ (see Figure 4(b)). Both vectors transform as $v^{\prime}_{1/2}=(R_{2\times 2}\oplus 1)v_{1/2}$ under rotation $R_{2\times 2}$, identical to the acceleration, and their cross product has the desirable property $\omega^{\prime}=v^{\prime}_{1}\times v^{\prime}_{2}=((R_{2\times 2}\oplus 1)v_{1})\times((R_{2\times 2}\oplus 1)v_{2})=\text{det}(R_{2\times 2}\oplus 1)(R_{2\times 2}\oplus 1)(v_{1}\times v_{2})=\text{det}(R_{2\times 2})(R_{2\times 2}\oplus 1)\omega$, using the standard cross-product property $(Ax)\times(Ay)=\text{det}(A)A(x\times y)$. The group action on $\omega$ exactly coincides with what was derived in Section 3. Our preprocessing step is summarized by the following mapping $\mathcal{F}:\mathbb{R}^{6}\rightarrow\mathbb{R}^{9}$ defined as:
@@
where we define $w_{1}=(-\omega_{y},\omega_{x},0)$ and $w_{2}=\omega\times w_{1}$. It can be checked that these vectors are indeed orthogonal and that $\omega=v_{1}\times v_{2}$. If $\omega_{x}=\omega_{y}=\bm{0}$, we use $w_{1}=a\times\omega$ and $w_{2}=\omega\times v_{1}$. Finally, if both $\omega_{x}=\omega_{y}=\bm{0}$ and $a\times\omega=\bm{0}$, we use $w_{1}=\omega\times(1,0,0)$ and $w_{2}=\omega\times v_{1}$. Applying an action $R_{2\times 2}$ on the input data results in
@@ 4.2 Equivariant Frame
Instead of enforcing the above equivariance constraint on every network layer, we instead employ a subequivariant network (see Figure 2(b)) to regress a canonical frame $G$ from the preprocessed input obtained from preprocessed IMU measurements $\{(a_{i},\omega_{i})\}_{i=1}^{n}$ when dealing with $SO_{\vec{g}}(3)$ and preprocessed IMU measurements $\{(a_{i},v_{1,i},v_{2,i})\}_{i=1}^{n}$, when dealing with $O_{\vec{g}}(3)$. The canonical frame $G$ is uniquely achieved by transforming the identity frame with $R_{G}\in SO(2)$ or $O(2)$, which means we equivalently learn the transformation $R_{G}$ when we learn the equivariant frame $G$. We first convert the IMU data into $n\times C_{0}^{s}$ rotation invariant scalar features, and $n\times 2\times C_{0}^{v}$ rotation equivariant vector features, and process them jointly, taking inspiration from [44], which shows that learning from invariant features alongside vector inputs can produce universally equivariant outputs. As vector features we select the $xy$-components of each input vector ($C_{0}^{v}=2$ for $SO(2)$ and $C_{0}^{v}=3$ for $O(2)$). Instead, as scalar features we select (i) the z-components of each vector, (ii) the norm of the $xy$-components of each vector, and (iii) the pairwise dot-product of the $xy$-components of each vector. For $SO(2)$ we have $C_{0}^{s}=2+2+1=5$, while for $O(2)$ we have $C_{0}^{s}=3+3+3=9$.
As seen in Figure 2(a), preprocessed IMU measurements $\{(a_{i},\omega_{i})\}_{i=1}^{n}$ are mapped into equivariant frame $G$ i.e. canonicalized into invariant features $\{(a^{\prime}_{i},\omega^{\prime}_{i})\}_{i=1}^{n}$. This mapping is $a^{\prime}_{i}=\rho_{a}(R_{G}^{T})a_{i}$, and similarly for other vectors. Note that we first perform the aforementioned preprocessing when dealing with $O_{\vec{g}}(3)$. We then regress the invariant displacement $d^{\prime}$ and covariance $\Sigma^{\prime}$ on these invariant features before mapping back into the original frame, using the same canonical frame. Specifically, the invariant covariance is remapped as $\text{vec}(\Sigma)=\rho_{\Sigma}(R_{G})\Sigma$. This results in equivariant displacement $d=\rho_{d}(R_{G})d^{\prime}$ and covariance $\text{vec}(\Sigma)=\rho_{\Sigma}(R_{G})\text{vec}(\Sigma^{\prime})$. This design enables seamless integration of our equivariant network while separating the effects of transformations, leveraging traditional networks to process intrinsic data.